OpenStudy (anonymous):

Find the derivative of the function. f(x) = (2x − 5)^4(x2 + x + 1)^5 ....the answer that I got was: f(x) = 20(2x-5)^3(2x)(x^2+x+1)^4(2x+1)

OpenStudy (anonymous):

Did you use product rule?

OpenStudy (anonymous):

I did but I think I went horribly wrong some where

OpenStudy (anonymous):

I can show you what i did

OpenStudy (anonymous):

Do you need the answer in any particular form? (fully simplified, fully factored, etc.)

OpenStudy (anonymous):

Fully simplified I'm guessing.

OpenStudy (anonymous):

hrm, that'll get annoying to fully expand that 5th power of a quadratic trinomial, but let's just get something correct for a derivative first.. What were your first 2 steps?

OpenStudy (anonymous):

f(x) = (2x − 5)4(x2 + x + 1)5 =4(2x-5)^3(2x)5(x^2+x+1)^4(2x+1) =4*5(2x-5)^3(2x)(x^2+x+1)^4(2x+1) =20(2x-5)^3(2x)(x^2+x+1)^4(2x+1)

OpenStudy (anonymous):

You're not adding the products correctly, and the derivative of 2x-5 is 2 not 2x.

OpenStudy (anonymous):

product rule is "first times the derivative of the second plus second times the derivative of the first."

OpenStudy (anonymous):

completely missed the 2x so the formula if like this? f'(x)g(x)+f(x)g'(x)

OpenStudy (anonymous):

Yes. \[Let \;s(x)=(2x-5)^4; \;t(x)=(x^2+x+1)^5. \] \[s'(x)=(2)(2x-5)^3; \; t'(x)=(5)(2x+1)(x^2+x+1)^4. \] \[\frac{d}{dx}s(x)t(x)=s(x)t'(x)+s'(x)t(x).\]

OpenStudy (anonymous):

Oops, forgot the (4) in the s'(x) there.

OpenStudy (anonymous):

Should read: \[s'(x)=(4)(2)(2x-5)^3\]

OpenStudy (anonymous):

That's the answer? really?

OpenStudy (anonymous):

It's the bits and pieces of the answer - you still need to put them all together and simplify.

OpenStudy (anonymous):

Remember that I took your f(x) and expressed it as f(x)=s(x)t(x) to showcase the product rule. s'(x) =/= f'(x).

OpenStudy (anonymous):

Now what do you have? (I recommend against trying to simplify it all the way down. )